Circular No. 8146

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

η CARINAE

M. F. Corcoran, Universities Space Research Association and Goddard Space Flight Center, reports that monitoring observations with the Rossi X-ray Timing Explorer (+ Proportional Counter Array) of η Car show an x-ray maximum on May 26 (∼ 5 days earlier than the predicted maximum; cf. IAUC 8137). Since that time, the x-ray flux has declined by 49 percent, and it is likely that the x-ray eclipse ingress has begun. A straight-line fit to the data since June 4 suggests that the x-ray minimum will occur on June 11 ± 2 days (formal error) if the decline remains linear. This is more than a month earlier than the previous predicted eclipse start (IAUC 8137). Additional ground-based observations are strongly encouraged.

SUPERNOVA 2003fa IN MCG +07-36-33 Further to IAUC 8142, T. Matheson et al. add that a spectrum (range

370–750 nm) of the nucleus of MCG +07-36-33, obtained by P. Berlind on June 6.37 UT, indicates that the galaxy has a recession velocity of 11800 km/s (so the NED value is in error) and thus is likely the host of SN 2003fa.

SUPERNOVAE 2003fb AND 2003fc M. Hamuy, Carnegie Observatories, reports that a spectrum (range

380–930 nm) of SN 2003fb (cf. IAUC 8143), obtained on June 7.33 UT with the Las Campanas Observatory Dupont 2.5-m telescope (+ WFCCD), confirms that this is a type-II supernova with well-developed Balmer lines exhibiting P-Cyg profiles. The minimum of the Hβ line yields an expansion velocity of −7054 km/s (adopting the NED host-galaxy recession velocity of 5259 km/s). A spectrum of SN 2003fc (cf. IAUC 8143), obtained on June 7.37, bears resemblance to that of the type-Ic SN 1994I obtained 4 days before maximum light. Superimposed narrow Hα emission from the host galaxy yields a redshift of 0.035.

R. J. Foley, K. M. Sandstrom, W. Li, and A. V. Filippenko, University of California, Berkeley, report that CCD spectra (range 320–1000 nm), obtained on June 7 UT with the Shane 3-m telescope at Lick Observatory, reveal that SN 2003fc is probably of type Ic near maximum brightness. There is a low-contrast, broad emission feature at the expected position of Hα, as is sometimes the case in type-Ic supernovae (e.g., Filippenko 1992, Ap.J. 384, L37). The recession speed of the host galaxy, measured from the narrow Hα emission of a superposed H II region, is ∼ 10400 km/s.

2003 June 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green